# Range, not Independence, Drives Modularity in Biological Inspired Representations

### v2

@@ 4.1 Linear RNNs
We study the optimal nonnegative, efficient, recurrent representations, ${\bm{z}}_{t}$, and show that whether the representations of the two frequencies within ${\bm{z}}_{t}$ modularises depends on their ratio. We prove and verify empirically that if one frequency is an integer multiple of the other the encodings mix, whereas if their ratio is irrational they should modularise. Further, we show empirically, and prove in limited settings, that rational non-harmonic ratios should modularise (Appendix E). The intuition for this result is much the same as the linear autoencoding setting: the natural notions of sources are the signals $(\cos(\omega_{1}t),\sin(\omega_{1}t),\cos(\omega_{2}t),\sin(\omega_{2}t))$. Using these sources, we must simply ask: does their support allow for a reduction in activity energy via mixing? In Figure 3a, we visualise the source support for the three prototypical relationships between $\omega_{1}$ and $\omega_{2}$: irrational, rational (but not harmonic), and harmonic. Results of neural network verifications are in Figure 3b (details Appendix I). In the irrational case, the source support is essentially rectangular, so the model modularises. In the harmonic case, large chunks of various corners are missing from the source support, so the model mixes. In the rational case, even though the source support is quite sparse, the corners are sufficiently present such that modularising is still optimal.
@@
Modularisation of grid cells. We now show that these spectral ideas can explain modules of grid cells (literature review: Appendix A. Grid cells are neurons in the mammalian entorhinal cortex that fire in a hexagonal lattice of positions [17]. They come in groups, called modules; grid cells within the same module have receptive fields (firing patterns) that are translated copies of the same lattice, and different modules are defined by their different lattice [48]. Current theories suggest that the grid cell system can be modelled as a RNN with activations built from linear combinations of frequencies [11], just like the linear RNNs considered here. Importantly, these grid cell theories use the same biological constraints as in this work and show that modules form because the optimal code contains non-harmonically related frequencies that are encoded in different neurons (Figure 3e top). This modularisation can now be theoretically justified in our framework, since non-harmonic frequencies are range-independent and so should be modularised (Figure 3e bottom).
@@
Figure 3: a) Source support visualisations: $\cos(\omega_{1}t)$ vs. $\cos(\omega_{2}t)$. b) The activity of the most mixed neuron in an optimal solution, the same neuron’s Fourier spectrum, and the population’s Fourier spectrum in the two task frequencies show clear modularisation in the irrational and rational cases, and clear mixing the in the harmonic case. c) Same as (b), but with nonlinear RNNs. d) Modularity of RNNs trained on 10 frequency pairs. e) Top: Two optimal grid cells from [11] in different modules. Bottom: The modularisation of these two lattices can be understood from the range-independence of their constituent frequencies, as shown in the joint distribution of the most significant frequency in each module. f) Plot of joint distribution of hidden state activity for two teacher RNNs neurons as we increase the dependence in three different ways. g) Trends of modularity scores of student RNN in three different cases qualitatively agree with our theory.
@@ 4.2 Nonlinear RNNs
Mixed sinusoidal regression for nonlinear RNNs. We train nonlinear ReLU RNNs with biologically inspired constraints to perform a frequency mixing task (details Appendix I). We provide a pulse input $P_{\omega}(t)=\mathbb{I}\left[\operatorname{mod}_{\omega}(t)=0\right]$ at two frequencies, and the network has to output the resulting “beats” and “carrier” signals:
@@
Modularisation in nonlinear teacher-student distillation. To test our predictions of when RNNs modularise, but in settings more realistic than pure frequencies, we generate training data trajectories from randomly initialised teacher RNNs with tanh activation function, and then train student RNNs (with a ReLU activation function) on these trajectories. The student’s representation is constrained to be nonnegative (via its ReLU) and has its activity and weights regularised (see Section I.2 for details). Using carefully chosen inputs at each timestep, we are able to precisely control the teacher RNN hidden state activity (i.e., the source distribution). This allows us to change correlations/statistical independence of the hidden states, while either maintaining or breaking range independence. We consider three settings (Figure 3e). First, when statistical and range independence are progressively broken (in orange). Second, where statistical, and to a lesser extent range, independence gets progressively broken (in green). Third, where only statistical independence gets progressively broken (in purple). We qualitatively observe, as our linear theory predicts, that students learns modular representations when range independence is maintained regardless of the statistical dependence of the teacher RNN hidden states (Figure 3f).
@@ 5 Modular or Mixed Codes for Space & Reward
We now apply our results to neuroscience to understand a puzzling difference in two seemingly similar recordings from entorhinal cortex. This brain area has been thought to contain precisely modular neurons, such as the grid cell code for self-position [17], object vector cells that fire at a particular displacement from objects [25], and heading direction cells [51]. Two recent papers examined the influence of rewarded locations on the grid cell code and find differing effects on the modularity of grid cells. Butler et al. [10] find that rewards rotate the grid cells, but preserve their pure spatial coding, while Boccara et al. [6] find that the grid cells warp towards the rewards, becoming mixed-selective to reward and space.
@@
Whittington et al. [56] study this discrepancy and point to the importance of the reward distribution in these two tasks: Boccara et al. [6] fix the positions of the possible rewards during one day, whereas Butler et al. [10] alternate the animals between periods of free-foraging for randomly placed rewards and periods of specific rewarded locations. However, the arguments of [56], which rely on source independence, are insufficient to explain these modularisation effects, as in neither case are reward and position independent; even in the experiments of Butler et al. [10] there are regions of space that are much more likely to be rewarded.
@@
Figure 4: a, b) We train linear RNNs to integrate a sequence of actions to report their current position and displacement from a reward. In some proportion of the rooms, rewards are randomly scattered; in the rest, they are in one of two fixed locations. c) Matching the theory, if some of the rooms are random then reward and position are range-independent, and their representation modularises, as in Stensola et al. [48] or Butler et al. [10]; if in all rooms the rewards occur in one of two fixed locations the representation mixes, with neurons encoding both reward and self position, as in Boccara et al. [6].
@@
However, with our improved understanding of modularisation, this makes sense. Despite Butler et al. [10] correlating reward and position, critically, they leave them range independent—all combinations of reward and position are possible. On the other hand, Boccara et al. [6] make certain combinations of reward and position impossible, making them not just correlated, but range correlated. As such, our theory matches this modularisation pattern. To test this we train a linear RNN with biological constraints (details Appendix J) to report both its self-position and displacement from a reward as it moves around an environment (Figure 4). We train RNNs on settings with different relationships between reward and position; for some RNNs the rewards are in fixed positions in every environment (range dependent and statistically dependent), for others reward and position are uniformly random in each room (range independent and statistically independent), and for a final group have both settings in different proportions (range independent but statistically dependent). As we vary the proportion of fixed rooms, we find the optimal representation stays modular, containing separate spatial and reward-vector cells as in Butler et al. [10], until all the sampled rooms have fixed reward positions, at which point neurons become mixed selective to reward and self-position, as in Boccara et al. [6]. As such, our theory now covers all known modularity results about grid cells.
@@
Missed sources and mixed selectivity. In contrast to the beautifully modular neurons in entorhinal cortex, recent work has highlighted neurons with mixed tuning to multiple navigational variables, such as position, heading direction, or speed [18]. We use our theory to highlight a potential artifact of this type of analysis. A purportedly mixed selective neuron may (in part) be purely selective for another unanalysed variable that is itself correlated with the measured spatial variables or behaviour. We highlight a simple example of this effect in Figure 5a, b. Imagine a mouse is keeping track of three objects as it moves in an environment (Figure 5a); we model this as a linear RNN with biological constraints that must report the displacement of all objects from itself (Figure 5b). We make the object positions range independent but correlated. As predicted by our theory, the RNN modularises these range-independent variables such that each neuron only encodes one object (Figure 5c). However, if an experimenter were only aware of two of the three objects, they would instead analyse the neural tuning with respect to those two objects. Due to the statistical dependencies between object positions, they would find mixed selective neurons that are in reality purely coding for the missing object (Figure 5d). We do not believe that all neurons in entorhinal cortex are modular; for example, there is clear evidence that neurons co-tuned to position and velocity are meaningfully mixed-selective, implementing the path-integration updates in the grid cell system [53]. Indeed, many of the neurons analysed in [18] likely are mixed selective. However, it is also seems likely that in more exploratory analyses such as these, mixed selectivity might often arise from missing encoded variables in the analysis.
@@ 6 Discussion
Limitations & future work. That said, there remain many aspects of our theory that need improvement. Deriving necessary (rather than sufficient) modularisation conditions for other norms and of multidimensional sources, extending to nonlinear settings, or to understand a less binarised notion of modularity (rather than perfectly modular or not) are attractive directions. Additionally, our model is a purposefully simple model, there are many biologically pertinent effects whose impact on modularity would be interesting to study, such as connection sparsity, anatomical constraints, noise, or requiring additional computational roles from the network. Finally, the theory needs testing beyond entorhinal cortex.
@@
Mixed selectivity. In neuroscience our findings add nuance to the ongoing debate over mixed vs. modular neural coding. Theories of nonlinear mixed selectivity have argued that, analogously to a nonlinear kernel, such schemes permit linear readouts to decode nonlinear functions of the sources [39]. This is likely a key part of the mixed selectivity found in some brain areas, like the cerebellum [31], mushroom body [1], and perhaps certain prefrontal or hippocampal representations [5, 7]. However, our theory raises the possibility for other explanations of both nonlinear and linear mixed selectivity that do not require tasks with nonlinear functions of the sources. First, purely from energy efficiency, our work analytically shows when range-dependent variables should be encoded in linearly mixed-selective representations, and empirically shows the same ideas apply to nonlinear mixed-selective representations Figure 2. Further, since in our theory even statistically dependent variables might be represented modularly, we show that this causes problems for analysis techniques that do not account for all the variables encoded in a population, leading to spurious mixed selectivity Section 5. Future work could usefully developed principled approaches to distinguishing these options.
@@
Disentangling what? In machine learning, our work bolsters the importance of range over statistical independence. We agree with Roth et al. [41], who argue that range independence is the more meaningful form of independence. Further, our conditions have qualitatively similarities to sufficient spread conditions from the matrix factorisation literature [49, 50], suggesting a deeper correspondence. It is intriguing that the constraints inspired by biology, which hold promise for disentangling [56], and have links to state-of-the-art disentangling methods [22, 23], are sensitive to precisely this form of independence.
Reproducibility: For code to reproduce the figures shown here please consult the following github repository: https://github.com/kylehkhsu/modular.
@@ Appendix A Related Works
Machine learning. The disentangled representation learning community has found inductive biases that lead networks to prefer modular or disentangled representations of nonlinear mixtures of independent variables. Examples include sparse temporal changes [29], smoothness or sparseness assumptions in the generative model [60, 20], and axis-wise latent quantisation [22, 23]. Relatedly, studies of tractable network models have identified a variety of structural aspects, in either the task or architecture, that lead to modularisation, including learning dynamics in gated linear networks [43], architectural constraints in linear networks [27, 46], compositional tasks in linear hypernetworks and nonlinear teacher-student frameworks [44, 33], or task-sparsity in linear autoencoders [16].
@@
Neuroscience. Theories prompted by the recognition of mixed-selectivity in biological neurons have argued that nonlinearly mixed selective codes might exist to enable a linear readout to flexibly decode any required categorisation [39], suggesting a generalisability-flexibility tradeoff between modular and nonlinear mixed encodings [5]. Modelling work has studied task-optimised network models of neural circuits, some of which have recovered mixed encodings [35]. However, other models trained on a wide variety of cognitive tasks have found that networks contain meaningfully modularised components [59, 12, 14]. Dubreuil et al. [13] study recurrent networks trained on similar cognitive tasks and find that, depending on the task structure, neurons in the trained networks can be grouped into populations defined by their connectivity patterns. To this developing body, our work and its predecessor [56] present understanding of when simple biological constraints would encourage encodings of many variables to be modularised into populations of single pure-tuned variables, or when mixed-selectivity would be preferred. Compared to [56] which only works when source variables are statistically independent, our work determines modularity for (almost) any dataset even when source variables are not independent. This is a dramatically more complete understanding of the boundary between mixed and modular codes that allows us to understand behaviour in both biological and artificial neurons that previously eluded explanation.
@@
Independent components analysis and matrix factorization. Our linear autoencoding problem (Theorem 2.1) bears some similarity to linear independent components analysis (ICA) and matrix factorization (MF). These assume data ${\bm{y}}_{i}$ is linearly generated from a set of constrained sources, ${\bm{y}}_{i}={\bm{A}}{\bm{s}}_{i}$, and build algorithms to recover ${\bm{s}}_{i}$ from ${\bm{y}}_{i}$. The literature studies different combinations of (1) assumptions on the sources ${\bm{s}}_{i}$, e.g. independence [4], nonnegativity[38], or boundedness [26]; (2) assumptions on the mixing matrix ${\bm{A}}$, e.g. nonnegative in nonnegative matrix factorisation and (3) algorithms, often deriving identifiability results: conditions that the data must satisfy for a given algorithm to succeed. Some work even derives biological implementations of such factorisation algorithms [37, 8]. In our work we study the conditions under which a particular biologically inspired representation modularises a set of bounded sources. We provide conditions on the sources under which the optimal representation is modular, morally similar to identifiability results (indeed, other identifiability results, like ours, sometimes take the form of sufficient spread conditions [49, 50]). We differ from previous work in the representation we study (which has only been studied before by [56]), the form of our identifiability result (which is both necessary and sufficient, has not been derived before, and assumes only that the sources are bounded), and these developments allows us to explain both recurrent modularisation and neural data.

### v4

@@ 4.1 Linear RNNs
We study the optimal nonnegative, efficient, recurrent representations, ${\bm{z}}_{t}$, and show that whether the representations of the two frequencies within ${\bm{z}}_{t}$ modularises depends on their ratio. We prove and verify empirically that if one frequency is an integer multiple of the other the encodings mix, whereas if their ratio is irrational they should modularise. Further, we show empirically, and prove in limited settings, that rational non-harmonic ratios should modularise (Appendix D). The intuition for this result is much the same as the linear autoencoding setting: the natural notions of sources are the signals $(\cos(\omega_{1}t),\sin(\omega_{1}t),\cos(\omega_{2}t),\sin(\omega_{2}t))$. Using these sources, we must simply ask: does their support allow for a reduction in activity energy via mixing? In Figure 3a, we visualise the source support for the three prototypical relationships between $\omega_{1}$ and $\omega_{2}$: irrational, rational (but not harmonic), and harmonic. Results of neural network verifications are in Figure 3b (details Appendix H). In the irrational case, the source support is essentially rectangular, so the model modularises. In the harmonic case, large chunks of various corners are missing from the source support, so the model mixes. In the rational case, even though the source support is quite sparse, the corners are sufficiently present such that modularising is still optimal.
@@
Modularisation of grid cells. We now show that these spectral ideas can explain modules of grid cells. Grid cells are neurons in the mammalian entorhinal cortex that fire in a hexagonal lattice of positions (Hafting et al., 2005). They come in groups, called modules; grid cells within the same module have receptive fields (firing patterns) that are translated copies of the same lattice, and different modules are defined by their different lattice (Stensola et al., 2012) (for further literature review see Appendix A). Current theories suggest that the grid cell system can be modelled as a RNN with activations built from linear combinations of frequencies (Dorrell et al., 2023), just like the linear RNNs considered here. Importantly, these grid cell theories use the same biological constraints as in this work and show that modules form because the optimal code contains non-harmonically related frequencies that are encoded in different neurons (Figure 3e top). This modularisation can now be theoretically justified in our framework, since non-harmonic frequencies are range-independent and so should be modularised (Figure 3e bottom).
@@
Figure 3: a) Source support visualisations: $\cos(\omega_{1}t)$ vs. $\cos(\omega_{2}t)$. b) The activity of the most mixed neuron in an optimal solution, the same neuron’s Fourier spectrum, and the population’s Fourier spectrum in the two task frequencies show clear modularisation in the irrational and rational cases, and clear mixing the in the harmonic case. c) Same as (b), but with nonlinear RNNs. d) Modularity of RNNs trained on 10 frequency pairs. e) Top: Two optimal grid cells from Dorrell et al. (2023) in different modules. Bottom: The modularisation of these two lattices can be understood from the range-independence of their constituent frequencies, as shown in the joint distribution of the most significant frequency in each module. f) Plot of joint distribution of hidden state activity for two teacher RNNs neurons as we increase the dependence in three different ways. g) Trends of modularity scores of student RNN in three different cases qualitatively agree with our theory.
@@ 4.2 Nonlinear RNNs
Mixed sinusoidal regression for nonlinear RNNs. To test how our ideas generalise beyond linear networks, we train nonlinear ReLU RNNs with biologically inspired constraints to perform a frequency mixing task (details Appendix H). We provide a pulse input $P_{\omega}(t)=\mathbb{I}\left[\operatorname{mod}_{\omega}(t)=0\right]$ at two frequencies, and the network has to output the resulting “beats” and “carrier” signals:
@@
Modularisation in nonlinear teacher-student distillation. To test our predictions of when RNNs modularise, but in settings more realistic than pure frequencies, we generate training data trajectories from randomly initialised teacher RNNs with tanh activation function, and then train student RNNs (with a ReLU activation function) on these trajectories. The student’s representation is constrained to be nonnegative (via its ReLU) and has its activity and weights regularised (see Section H.2 for details). Using carefully chosen inputs at each timestep, we are able to precisely control the teacher RNN hidden state activity (i.e., the source distribution). This allows us to change correlations/statistical independence of the hidden states, while either maintaining or breaking range independence. We consider three settings (Figure 3f). First, when statistical and range independence are progressively broken (in orange). Second, where statistical, and to a lesser extent range, independence gets progressively broken (in green). Third, where only statistical independence gets progressively broken (in purple). We observe that, in line with our theory, preserving the corner points preserves modularity (Figure 3g purple), while removing them breaks it (orange). Further, spasifying the corners breaks modularity, but more slowly than removing corners (green).
@@ 5 Modular or Mixed Codes for Space & Reward
We now apply our results to neuroscience to understand a puzzling difference in two seemingly similar recordings from entorhinal cortex. This brain area has been thought to contain precisely modular neurons, such as the grid cell code for self-position (Hafting et al., 2005), object vector cells that fire at a particular displacement from objects (Høydal et al., 2019), and heading direction cells (Taube, 2007). Two recent papers examined the influence of rewarded locations on the grid cell code and find differing effects on the modularity of grid cells. Butler et al. (2019) find that rewards rotate the grid cells, but preserve their pure spatial coding, while Boccara et al. (2019) find that the grid cells warp towards the rewards, becoming mixed-selective to reward and space.
@@
Whittington et al. (2023) study this discrepancy and point to the importance of the reward distribution in these two tasks: Boccara et al. (2019) fix the positions of the possible rewards during one day, whereas Butler et al. (2019) alternate the animals between periods of free-foraging for randomly placed rewards and periods of specific rewarded locations. However, the arguments of Whittington et al. (2023), which rely on source independence, are insufficient to explain these modularisation effects, as in neither case are reward and position independent; even in the experiments of Butler et al. (2019) there are regions of space that are much more likely to be rewarded.
@@
Figure 4: a, b) We train linear RNNs to integrate a sequence of actions to report their current position and displacement from a reward. In some proportion of the rooms, rewards are randomly scattered; in the rest, they are in one of two fixed locations. c) Matching the theory, if some of the rooms are random then reward and position are range-independent, and their representation modularises, as in Stensola et al. (2012) or Butler et al. (2019); if in all rooms the rewards occur in one of two fixed locations the representation mixes, with neurons encoding both reward and self position, as in Boccara et al. (2019).
@@
However, with our improved understanding of modularisation, this makes sense. Despite Butler et al. (2019) correlating reward and position, critically, they leave them range independent—all combinations of reward and position are possible. On the other hand, Boccara et al. (2019) make certain combinations of reward and position impossible, making them not just correlated, but range correlated. As such, our theory matches this modularisation pattern. To test this we train a linear RNN with biological constraints (details Section H.3) to report both its self-position and displacement from a reward as it moves around an environment (Figure 4). We train RNNs on settings with different relationships between reward and position; for some RNNs the rewards are in fixed positions in every environment (range dependent and statistically dependent), for others reward and position are uniformly random in each room (range independent and statistically independent), and a final group experiences both settings in different proportions (range independent but statistically dependent). As we vary the proportion of fixed rooms, we find the optimal representation stays modular, containing separate spatial and reward-vector cells as in Butler et al. (2019), until all the sampled rooms have fixed reward positions, at which point neurons become mixed selective to reward and self-position, as in Boccara et al. (2019). As such, our theory now covers all known grid cell modularity results.
@@
Missed sources and mixed selectivity. In contrast to the beautifully modular neurons in entorhinal cortex, recent work has highlighted neurons with mixed tuning to multiple navigational variables, such as position, heading direction, or speed (Hardcastle et al., 2017). We use our theory to highlight a potential artifact of this type of analysis. A purportedly mixed selective neuron may (in part) be purely selective for another unanalysed variable that is itself correlated with the measured spatial variables or behaviour. We highlight a simple example of this effect in Figure 5. Imagine a mouse is keeping track of three objects as it moves in an environment (Figure 5a); we model this as a linear RNN with biological constraints that must report the displacement of all objects from itself. We make the object positions range independent but correlated. As predicted by our theory, the RNN modularises these range-independent variables such that each neuron only encodes one object (Figure 5c). However, if an experimenter were only aware of two of the three objects, they would instead analyse the neural tuning with respect to those two objects, without being able to condition on the third. Due to the statistical dependencies between object positions, they would find mixed selective neurons that are in reality purely coding for the missing object (Figure 5d). We do not believe that all neurons in entorhinal cortex are modular; for example, there is clear evidence that neurons co-tuned to position and velocity are meaningfully mixed-selective, implementing the path-integration updates in the grid cell system (Vollan et al., 2025). Indeed, many of the neurons analysed in Hardcastle et al. (2017) likely are mixed selective. However, it is also seems likely that in more exploratory analyses such as these, mixed selectivity might often arise from missing encoded variables in the analysis.
@@ 6 Discussion
Limitations & future work. That said, there remain many aspects of our theory that need improvement. Deriving necessary (rather than just sufficient) modularisation conditions for other norms and of multidimensional sources, extending to nonlinear settings, or to understand a less binarised notion of modularity (rather than perfectly modular or not), or developing a theory of linearly mixed sources as in matrix factorsiation work, are all attractive directions. Further, we assume perfect reconstruction. This leads us to some of our suprising conclusions, like how it is the support of the distribution that matters, even if some parts of the support have very low probability. In Appendix C we show that allowing the representation to tradeoff reconstruction accuracy with the other losses leads it to ignore very low probability points. Studying this tradeoff further would be interesting. Biologically, beyond the clear need for experimental validation, there are many pertinent effects whose impact on modularity would be interesting to study, such as connection sparsity, anatomical constraints, noise, or requiring additional computational roles from the network.
@@
Mixed selectivity. In neuroscience our findings add nuance to the ongoing debate over mixed vs. modular neural coding. Theories of nonlinear mixed selectivity have argued that, analogously to a nonlinear kernel, such schemes permit linear readouts to decode nonlinear functions of the sources (Rigotti et al., 2013). This is likely a key part of the mixed selectivity found in some brain areas, like the cerebellum (Lanore et al., 2021), mushroom body (Aso et al., 2014), and perhaps certain prefrontal or hippocampal representations (Bernardi et al., 2020; Boyle et al., 2024). However, our theory raises the possibility for other explanations of both nonlinear and linear mixed selectivity that do not require tasks with nonlinear functions of the sources. First, purely from energy efficiency, our work analytically shows when range-dependent variables should be encoded in linearly mixed-selective representations, and empirically shows the same ideas apply to nonlinear mixed-selective representations, Figure 2. Further, since in our theory even statistically dependent variables might be represented modularly, we show that this causes problems for analysis techniques that do not account for all the variables encoded in a population, leading to spurious mixed selectivity, Figure 5. Future work could usefully develop principled approaches to distinguishing these options.
@@
Disentangling what? We agree with Roth et al. (2023), who argue that range independence is the more meaningful form of independence for modularising variables. Further, the similarity between our conditions and those from the matrix factorisation literature (e.g. (Tatli and Erdogan, 2021a; b)), point to repeated instances of similar ideas. It is intriguing that constraints inspired by biology, which hold promise for disentangling (Whittington et al., 2023), and have links to state-of-the-art disentangling methods (Hsu et al., 2023; 2024), are sensitive to precisely this form of independence.
Reproducibility Statement Our theoretical work is fully detailed in the appendices, and code to reproduce our empirical work can be found at https://github.com/kylehkhsu/modular.
Acknowledgements The authors thank Basile Confavreux, Pierre Glaser, & Bariscan Bozkurt for conversations about proof techniques.
We thank the following funding sources: Gatsby Charitable Foundation to W.D., J.H.L. & P.E.L.; Sequoia Capital Stanford Graduate Fellowship and NSERC Postgraduate Scholarship – Doctoral to K.H.; Wellcome Trust (110114/Z/15/Z) to P.E.L.; Wellcome Trust (219627/Z/19/Z) to J.H.L.; Sir Henry Wellcome Postdoctoral Fellowship (222817/Z/21/Z) to JCRW.; Wellcome Principal Research Fellowship (219525/Z/19/Z), Wellcome Collaborator award (214314/Z/18/Z), and Jean-François and Marie-Laure de Clermont-Tonnerre Foundation award (JSMF220020372) to TEJB.; the Wellcome Centre for Integrative Neuroimaging is supported by core funding from the Wellcome Trust (203139/Z/16/Z).
@@ Appendix A Related Works
Machine learning. The disentangled representation learning community has found inductive biases that lead networks to prefer modular or disentangled representations of nonlinear mixtures of independent variables. Examples include sparse temporal changes [Klindt et al., 2020], smoothness or sparseness assumptions in the generative model [Zheng et al., 2022, Horan et al., 2021], and axis-wise latent quantisation [Hsu et al., 2023, 2024]. Relatedly, studies of tractable network models have identified a variety of structural aspects, in either the task or architecture, that lead to modularisation, including learning dynamics in gated linear networks [Saxe et al., 2022], architectural constraints in linear networks [Jarvis et al., 2023, Shi et al., 2022], compositional tasks in linear hypernetworks and nonlinear teacher-student frameworks [Schug et al., 2024, Lee et al., 2024], or task-sparsity in linear autoencoders [Elhage et al., 2022].
@@
Neuroscience. Theories prompted by the recognition of mixed-selectivity in biological neurons have argued that nonlinearly mixed selective codes might exist to enable a linear readout to flexibly decode any required categorisation [Rigotti et al., 2013], suggesting a generalisability-flexibility tradeoff between modular and nonlinear mixed encodings [Bernardi et al., 2020]. Modelling work has studied task-optimised network models of neural circuits, some of which have recovered mixed encodings [Nayebi et al., 2021]. However, other models trained on a wide variety of cognitive tasks have found that networks contain meaningfully modularised components [Yang et al., 2019, Driscoll et al., 2024, Duncker et al., 2020]. Dubreuil et al. [2022] study recurrent networks trained on similar cognitive tasks and find that, depending on the task structure, neurons in the trained networks can be grouped into populations defined by their connectivity patterns. To this developing body our work and its predecessor [Whittington et al., 2023] add understanding of when simple biological constraints would encourage encodings of many variables to be modularised into populations of single pure-tuned variables, or when mixed-selectivity would be preferred. Compared to [Whittington et al., 2023] which only works when source variables are statistically independent, our work determines whether the optimal representation is modular for any dataset even when source variables are not independent. This complete understanding of the boundary between mixed and modular codes is what allows us to understand behaviour in both biological and artificial neurons that previously eluded explanation.
@@
Independent components analysis and matrix factorization. Our linear autoencoding problem (Theorem 2.1) bears some similarity to linear independent components analysis (ICA) and matrix factorization (MF). These assume data ${\bm{y}}_{i}$ is linearly generated from a set of constrained sources, ${\bm{y}}_{i}={\bm{A}}{\bm{s}}_{i}$, and build algorithms to recover ${\bm{s}}_{i}$ from ${\bm{y}}_{i}$. The literature studies different combinations of (1) assumptions on the sources ${\bm{s}}_{i}$, e.g. independence [Bell and Sejnowski, 1995], nonnegativity[Plumbley, 2003], or boundedness Inan and Erdogan [2014]; (2) assumptions on the mixing matrix ${\bm{A}}$, e.g. nonnegative in nonnegative matrix factorisation and (3) algorithms, often deriving identifiability results: conditions that the data must satisfy for a given algorithm to succeed. Some work even derives biological implementations of such factorisation algorithms [Pehlevan et al., 2017, Bozkurt et al., 2022]. In our work we study the conditions under which a particular biologically inspired representation modularises a set of bounded sources. We provide conditions on the sources under which the optimal representation is modular, morally similar to identifiability results (indeed, other identifiability results, like ours, sometimes take the form of sufficient spread conditions [Tatli and Erdogan, 2021a, b]). We differ from previous work in the representation we study (which has only been studied before by Whittington et al. [2023]), the form of our identifiability result (which is both necessary and sufficient, and has not been derived before), and how we apply these ideas, for example to neural data.